### v1

@@ 3.4 A Brief Analysis of Model Complexity
In Section 2.3, we claim that our Hypernetworks-based method can work better than conditional embedding based methods like AirNet [20] and PromptIR [24], so we will give a simple proof for this. In the context of all-in-one image restoration, we aim to find a function $f:\mathbb{R}^{3HW}\xrightarrow{}\mathbb{R}^{3HW}$ to map a degraded image $\mathbf{X}$ to a restored image $\hat{\mathbf{X}}$, minimizing the distance between $\hat{\mathbf{X}}$ and the ground truth $\mathbf{Y}=y(\mathbf{X})$. (For the sake of discussion, the following tensors are generally regarded as flattened one-dimensional vectors) $(a)$ Mainstream methods today typically utilize a neural network $g(\mathbf{X},\mathbf{I(\mathbf{X})})$ to learn the mapping, where $\mathbf{I(\mathbf{X})}\in\mathbb{R}^{m}$ stands for the degradation information (degradation embedding) from the input image, such as the prompt in PromptIR [24]. These methods directly send $\mathbf{I}(\mathbf{X})$ together with $\mathbf{X}$ into the network to get $\hat{\mathbf{X}}$. $(b)$ For our HAIR, the neural network can be formulated as $h(\mathbf{X};\theta(\mathbf{I(\mathbf{X})}))$, where $\theta$ is a function that maps $\mathbf{I(\mathbf{X})}$ into the parameters of $h$. We define the distance between 2 functions as follows:
$d(g,y)=\min_{g}\max_{\mathbf{X}}\|g(\mathbf{X},\mathbf{I}(\mathbf{X}))-y(\mathbf{X})\|_{\infty}$ (8)
Given that vector functions can be complex, we define scalar function $f^{i}:\mathbb{R}^{M}\xrightarrow{}\mathbb{R},i=1,2,3\cdots M.$ of a vector function $f:\mathbb{R}^{M}\xrightarrow{}\mathbb{R}^{M}$. $i$ means $f^{i}$ produce the $i$-th value of $f$, then we have:
@@
$=\min_{g}\max_{\mathbf{X}}|g^{k(g)}(\mathbf{X})-y^{k(g)}(\mathbf{X})|$
@@
Intuitively, since $y^{i}$ maps the degraded image to a single pixel of the clean image, a little disturbance on the degraded image should only bring a little difference to the single pixel, so $y^{i}$ should be smooth and differentiable or at least continue. What’s more, the domain and range of $y$ are restricted to $[0,1]^{3HW}$, with most of the pixels far more less than 1, the Sobolev norm of $y_{i}$ can be generally less than 1. So this assumption generally holds.
Assumption 2
@@
In our context, the various $g$ actually stands for the same network with different weights during training. No matter how the parameters updates, the most "difficult" degraded image genrally remains the same one. (e.g., the one with very severe degradation) And within this "difficult" image, the most "difficult" pixel should also remains the same. (e.g., the pixel that varies dramatically from its clean version.) In this way, we simplify the vector function $g$ to scalar function $g^{k(g)}$. With these 2 assumptions, incorporating the Theorem 2, Theorem 3 Theorem 4 and all the assumptions in [32], we have the following conclusions:
Theorem 1
For conditional embedding based all-in-one image restoration methods $g(\mathbf{X},\mathbf{I(\mathbf{X})})$ like PromptIR [24], to achieve error $d(g,y)\leq\epsilon$, the complexity (number of parameters) of the corresponding model is at least: $N_{g}=\Omega(\epsilon^{-3HW-m})$
Theorem 2
For Hypernetworks based methods $h(\mathbf{X};\theta(\mathbf{I}(\mathbf{X})))$ like our HAIR, to achieve error $d(h,y)\leq\epsilon$, the complexity of the corresponding model is: $N_{\theta}=O(\epsilon^{-3HW/r})$
See paper [32] for detailed premises and proof of Theorem 1, 2. In this way, we show that to achieve the same error, HAIR will cost less parameters than conditional embedding based methods.
@@ 4 Experiments
We conduct experiments by rigorously adhering to the methodologies established in prior research on general image restoration [25, 24], under two distinct experimental configurations: (a) All-in-One and (b) Single-task. Under the All-in-One configuration, a singular model is trained to handle multiple degradation types, encompassing both three and five unique degradations. In contrast, the Single-task configuration involves the training of individual models dedicated to specific restoration tasks.
@@ 4.1 Experimenting Preparation
Implementation Details. Our HAIR framework is designed to be end-to-end trainable, eliminating the need for pre-training any of its components. The architecture comprises a 4-level encoder-decoder structure, with each level equipped with a varying number of Transformer blocks. Specifically, the number of blocks increases progressively from level-1 to level-4, following the sequence [4, 6, 6, 8]. HyperTrans Blocks are employed throughout level-4 and the decoding stages of levels 1-3. Additionally, the Weights Box parameter $N$ is set according to the sequence [5, 7, 7, 9] for each respective level. In the all-in-one setting, the model is trained with a batch size of 32, while in the single-task setting, a batch size of 8 is used. The network’s optimization is guided by an L1 loss function, employing the AdamW optimizer [48] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $42e-$ 4 for the initial 150 epochs, and we change it to $52e-$ 5 for the final 10 epochs. To enhance the training data, input patches of size 128 x 128 are utilized, with random horizontal and vertical flips applied to the images to augment the dataset.
Datasets. We follow the approaches in previous studies [20, 24, 25] for our All-in-One and single-task experiments, using the following datasets: For single-task image denoising, we use the BSD400 [49] and WED [50] datasets, adding Gaussian noise at levels $\sigma\in\{15,25,50\}$ to generate training images. Testing is performed on the BSD68 [51] and Urban100 [52] datasets. For deraining, we employ the Rain100L dataset [53]. Dehazing experiments utilize the SOTS dataset [54]. Deblurring and low-light enhancement tasks use the GoPro [55] and LOL-v1 [56] datasets respectively. To create a comprehensive model for all tasks, we combine these datasets and train in a setting that covers three or five types of degradations. For single-task models, training is conducted for the same duration on the respective dataset.
Table 1: Comparison to state-of-the-art on three degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best and second best performances are highlighted. Our method sets a new state-of-the-art on average across all benchmarks while being significantly more efficient than prior work. ‘-’ represents unreported results.
@@ 4.2 Results
All-in-One: Three Degradations. We conducted a comparative evaluation of our All-in-One image restoration framework against several specialized state-of-the-art methods, including BRDNet [57], LPNet [58], FDGAN [59], DL [23], MPRNet [21], AirNet [20], PromptIR [24] and DaAIR [26]. Each of these methods was trained to handle the tasks of dehazing, deraining, and denoising. As illustrated in Table 1, our approach excels as the leading contender, offering a significant average improvement of $0.64$ dB over PromptIR [24]. Our method notably outperforms existing benchmarks on BSD68 and Rain100L datasets, setting a new precedent by exceeding the performance of the previous best methods by margins of $0.19$ dB and $1.49$ dB respectively.
All-in-One: Five Degradations. Following recent studies [25, 26], we have extended our approach to include deblurring and low-light image enhancement, thereby broadening our three-degradation framework to a more complex All-in-One setting. As demonstrated in Table 2, our method excels by learning specialized models for each degradation type while effectively capturing the shared characteristics between tasks. It achieves an average improvement of $0.6$ dB over PromptIR [24] and $0.13$ dB over DaAIR [26], establishing a new benchmark for state-of-the-art (SOTA) performance across all five benchmarks. Additionally, we compared our method with general image restoration models trained in the same All-in-One setting. Notably, our method significantly outperforms our baseline, Restormer [12], in various tasks. This comparison highlights the strong capability of our method as a versatile plug-in-and-play module, enhancing the performance of existing restoration architectures with minimal integration complexity.
@@
DaAIR (ours) 37.78 .982
HAIR (ours) 39.00 .985
@@
HAIR (ours) 34.93 .953 32.70 .931 29.65 .885
Visual Results The visual results captured under three degradation scenarios are presented in Fig. 4. These results clearly demonstrate the superiority of our method in terms of visual quality: In the denoising task at $\sigma=25$, HAIR retains more fine details compared to other methods; in the deraining task, HAIR effectively removes all rain streaks, whereas the compared methods exhibit noticeable remnants; in the dehazing task, HAIR may not closely resemble the Ground Truth, yet it provides the most visually pleasing outcome and even clears the original haze present in the Ground Truth. Collectively, these observations confirm the efficacy of our approach in enhancing image quality across different degradation conditions.
@@ 4.3 Ablation Studies
Effectiveness of Classifier and Hyper Selecting Net (HSN) As detailed in Table 4, we study the impact of the proposed modules in four settings: (a) Here, the model aligns with the Restormer architecture [12]. (b) In this configuration, the Global Information Vector (GIV) is directly used as conditional embeddings for the Transformer Blocks in the Decoder, rather than for weight generation. (c) The GIVs are designated as independently trainable parameters, each randomly initialized, with the Decoder’s Transformer Blocks having their own distinct GIVs. (d) This setup incorporates both components. The results demonstrate the indispensability of both components. With the addition of only 3M parameters and no change to the logical structure, HAIR outperforms Restormer by 1.7 dB in PSNR, underscoring the simplicity and effectiveness of our proposed approach.
@@
HAIR for Other Baselines. We have previously posited that HAIR is readily integrable with any image-to-image network architecture. To substantiate this claim, we have implemented HAIR on various established baselines. As depicted in Table Table 6, we selected three efficacious image restoration models—Transweather [1], AirNet [20], and Restormer [12]—for integration with HAIR. Specifically, we have integrated the Classifier at the network’s midpoint for each method and transitioned the subsequent layers to Hypernetworks-based modules. The outcomes shows that our method outperforms PromptIR as a plug-in-and-play module, underscoring the universality and potency of our approach.
@@
Method Architecture Params Dehazing Deraining Denoising Deblurring Low-Light Average
@@ 5 Conclusion
This paper introduces HAIR, a novel plug-and-play Hypernetworks-based module capable of being easily integrated and adaptively generating parameters for different networks based on the input image. Our method comprises two main components: the Classifier and the Hyper Selecting Net (HSN). Specifically, the Classifier is a simple image classification network with Global Average Pooling, designed to produce a Global Information Vector (GIV) that encapsulates the global information from the input image. The HSN functions as a Fully-connected Neural Network, receiving the GIV and outputting parameters for the corresponding modules. Extensive experiments indicate that HAIR can significantly enhance the performance of various image restoration architectures at a low cost, without necessitating any changes to their logical structures. By incorporating HAIR into the widely recognized Restormer architecture, we have achieved State-Of-The-Art performance on a range of image restoration tasks. The potential for further exploitation of data-conditioned Hypernetworks in image-to-image tasks is substantial, considering their robust adaptability to diverse input images based on their content.

### v2

@@ 3.4 A Brief Analysis of Model Complexity
In Section 2.3, we claim that our Hypernetworks-based method can work better than conditional embedding-based methods like AirNet [20] and PromptIR [24]. This section will give a simple proof for this point. In the context of all-in-one image restoration, we aim to find a function $f:\mathbb{R}^{3HW}\xrightarrow{}\mathbb{R}^{3HW}$ to map a degraded image $\mathbf{X}$ to a restored image $\hat{\mathbf{X}}$ by minimizing the distance between $\hat{\mathbf{X}}$ and the ground truth $\mathbf{Y}=y(\mathbf{X})$. [^1] For the sake of discussion, the following tensors are generally regarded as flattened one-dimensional vectors. Mainstream methods typically utilize a neural network $g(\mathbf{X},e(\mathbf{X}))$ to learn the mapping, where $e(\mathbf{X})\in\mathbb{R}^{k}$ stands for the degradation information (degradation embedding) from the input image, such as the prompt in PromptIR [24]. These methods directly send $e(\mathbf{X})$ together with $\mathbf{X}$ into the network to get $\hat{\mathbf{X}}$. For our HAIR, the neural network can be formulated as $h(\mathbf{X};\theta(e(\mathbf{X})))$, where $\theta$ is a function that maps ${e(\mathbf{X})}$ into the parameters of $h$. We define the distance between the two functions as:
$d(g,y)=\min_{g}\max_{\mathbf{X}}\|g(\mathbf{X},e(\mathbf{X}))-y(\mathbf{X})\|_{\infty}.$ (8)
Given that vector functions can be complex, we define a scalar function $f^{i}:\mathbb{R}^{M}\xrightarrow{}\mathbb{R},i=1,2,3\cdots M.$ of a vector function $f:\mathbb{R}^{M}\xrightarrow{}\mathbb{R}^{M}$. $i$ means $f^{i}$ produce the $i$-th value of $f$, then we have:
@@
$=\min_{g}\max_{\mathbf{X}}|g^{k(g)}(\mathbf{X})-y^{k(g)}(\mathbf{X})|.$
@@
Intuitively, since $y^{i}$ maps the degraded image to a single pixel of the clean image, a little disturbance on the degraded image should only bring a little difference to the single pixel, so $y^{i}$ should be smooth and differentiable or at least continue. What’s more, the domain and range of $y$ are restricted to $[0,1]^{3HW}$, with most of the pixels far less than 1, the Sobolev norm of $y_{i}$ can be generally less than 1. So this assumption generally holds.
Assumption 2.
@@
In our context, the various $g$ represents the same network with different weights during training. No matter how the parameters update, the most "difficult" degraded image generally remains the same one, e.g., the one with very severe degradation. Within this "difficult" image, the most "difficult" pixel should also remain the same, e.g., the pixel that varies dramatically from its clean version. In this way, we simplify the vector function $g$ to scalar function $g^{k(g)}$. With these two assumptions, as well as incorporating Theorem 2-4 and all the assumptions in [32], we can obtain the following conclusion.
Theorem 1.
For conditional embedding-based all-in-one image restoration methods $g(\mathbf{X},e(\mathbf{X}))$ like PromptIR [24], to achieve error $d(g,y)\leq\epsilon$, the complexity (number of parameters) of the corresponding model is at least $N_{g}=\Omega(\epsilon^{-3HW-k})$.
Theorem 2.
For Hypernetworks based methods $h(\mathbf{X};\theta(e(\mathbf{X})))$ like our HAIR, to achieve error $d(h,y)\leq\epsilon$, the complexity of the corresponding model is $N_{\theta}=O(\epsilon^{-3HW/r})$.
Please refer to paper [32] and Section 6.1 for detailed premises and proof of Theorem 1 and 2. The two Theorems demonstrate that to achieve the same error $\epsilon$, our proposed HAIR requires fewer parameters than conditional embedding-based methods. (As long as $\epsilon$ is small enough.)
@@ 4 Experiments
We conduct experiments by rigorously adhering to the methodologies established in prior research on general image restoration [25, 24], under two distinct experimental configurations: (a) All-in-One and (b) Single-task. In the All-in-One configuration, a singular model is trained to handle multiple types of degradation, encompassing both three and five unique degradations. In contrast, the Single-task configuration involves the training of individual models dedicated to specific restoration tasks.
@@ 4.1 Experimenting Preparation
Implementation Details. Our Res-HAIR framework is designed to be end-to-end trainable, eliminating the need for pre-training any of its components. The architecture comprises a 4-level encoder-decoder structure, with each level equipped with a varying number of Transformer blocks. Specifically, the number of blocks increases progressively from level-1 to level-4, following the sequence [4, 6, 6, 8]. HyperTrans Blocks are employed throughout level-4 and the decoding stages of levels 1-3. Additionally, the Weights Box parameter $N$ is set according to the sequence [5, 7, 7, 9] for each respective level. In the all-in-one setting, the model is trained with a batch size of 32, while in the single-task setting, a batch size of 8 is used. The optimization of the network is guided by an L1 loss function, employing the AdamW optimizer [48] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $42e-$ 4 for the initial 150 epochs, and we change it to $52e-$ 5 for the final 10 epochs. To enhance the training data, input patches of size 128 $\times$ 128 are utilized, with random horizontal and vertical flips applied to the images to augment the dataset.
Datasets. We follow the approaches in previous studies [20, 24, 25] for our All-in-One and single-task experiments, using the following datasets. For single-task image denoising, we use the BSD400 [49] and WED [50] datasets, adding Gaussian noise at levels $\sigma\in\{15,25,50\}$ to generate training images. Testing is performed on the BSD68 [51] and Urban100 [52] datasets. For deraining, we employ the Rain100L dataset [53]. Dehazing experiments utilize the SOTS dataset [54]. Deblurring and low-light enhancement tasks use the GoPro [55] and LOL-v1 [56] datasets respectively. To create a comprehensive model for all tasks, we combine these datasets and train them in a setting that covers three or five types of degradations. For single-task models, training is conducted for the same duration on the respective dataset.
Table 1: Comparison to state-of-the-art on three degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best and second best performances are highlighted. Our method sets a new state-of-the-art on average across all benchmarks while being significantly more efficient than previous work. ‘-’ represents unreported results.
@@ 4.2 Results
All-in-One: Three degradations. We conducted a comparative evaluation of our All-in-One image restoration framework against several state-of-the-art specialized methods, including BRDNet [57], LPNet [58], FDGAN [59], DL [23], MPRNet [21], AirNet [20], PromptIR [24], and DaAIR [26]. Each of these methods was trained to handle the tasks of dehazing, deraining, and denoising. As illustrated in Table 1, our approach excels as the leading contender, offering a significant average improvement of $0.64$ dB over PromptIR [24]. Our method notably outperforms existing benchmarks on BSD68 and Rain100L datasets, setting a new precedent by exceeding the performance of the previous best methods by margins of $0.19$ dB and $1.49$ dB respectively.
All-in-One: Five Degradations. Following recent studies [25, 26], we have extended our approach to include deblurring and low-light image enhancement, thereby broadening our three-degradation framework to a more complex All-in-One setting. As demonstrated in Table 2, our method excels by learning specialized models for each degradation type while effectively capturing the shared characteristics between tasks. It achieves an average improvement of $0.6$ dB over PromptIR [24] and $0.13$ dB over DaAIR [26], establishing a new benchmark for state-of-the-art (SOTA) performance across all five benchmarks. Additionally, we compared our method with general image restoration models trained in the same All-in-One setting. Notably, our method significantly outperforms our baseline (i.e. Restormer [12]) in various tasks. This comparison highlights the strong capability of our method as a versatile plug-in-and-play module, enhancing the performance of existing restoration architectures with minimal integration complexity.
@@
DaAIR [26] 37.78 .982
Res-HAIR (ours) 39.00 .985
@@
Res-HAIR (ours) 34.93 .953 32.70 .931 29.65 .885
Visual Results The visual results captured under three degradation scenarios are presented in Fig. 4. These results clearly demonstrate the superiority of our method in terms of visual quality: In the denoising task at $\sigma=25$, Res-HAIR retains more fine details compared to other methods; in the deraining task, Res-HAIR effectively removes all rain streaks, whereas the compared methods exhibit noticeable remnants; in the dehazing task, Res-HAIR may not closely resemble the Ground Truth, yet it provides the most visually pleasing outcome and even clears the original haze present in the Ground Truth. Collectively, these observations confirm the efficacy of our approach in enhancing image quality across different degradation conditions.
@@ 4.3 Ablation Studies
Effectiveness of Classifier and Hyper Selecting Net (HSN) As detailed in Table 4, we study the impact of the proposed modules in four settings: (a) Here, the model aligns with the Restormer architecture [12]. (b) In this configuration, the Global Information Vector (GIV) is directly used as conditional embeddings for the Transformer Blocks in the Decoder, rather than for weight generation. (c) The GIVs are designated as independently trainable parameters, each randomly initialized, with the Decoder’s Transformer Blocks having their own distinct GIVs. (d) This setup incorporates both components. The results demonstrate the indispensability of both components. With the addition of only 3M parameters and no change to the logical structure, Res-HAIR outperforms Restormer by 1.7 dB in PSNR, underscoring the simplicity and effectiveness of our proposed approach.
@@
HAIR for Other Baselines. We have previously posited that HAIR is essentially a plug-in-and-play module which is readily integrable with any image-to-image network architecture. To substantiate this claim, we have implemented HAIR on various established baselines. As depicted in Table Table 6, we selected three efficacious image restoration models—Transweather [1], AirNet [20], and Restormer [12]—for integration with HAIR. Specifically, we have integrated the Classifier at the network’s midpoint for each method and transitioned the subsequent layers to Hypernetworks-based modules. The outcomes show that our method outperforms PromptIR as a plug-in-and-play module, underscoring the universality and potency of our approach.
@@
Method Main Operation Params Dehazing Deraining Denoising Deblurring Low-Light Average
@@ 5 Conclusion
This paper introduces HAIR, a novel plug-and-play Hypernetworks-based module capable of being easily integrated and adaptively generating parameters for different networks based on the input image. Our method comprises two main components: the Classifier and the Hyper Selecting Net (HSN). Specifically, the Classifier is a simple image classification network with Global Average Pooling, designed to produce a Global Information Vector (GIV) that encapsulates the global information from the input image. The HSN functions as a Fully-connected Neural Network, receiving the GIV and outputting parameters for the corresponding modules. Extensive experiments indicate that HAIR can significantly enhance the performance of various image restoration architectures at a low cost without necessitating any changes to their logical structures. By incorporating HAIR into the widely recognized Restormer architecture, we have achieved State-Of-The-Art performance on a range of image restoration tasks. The potential for further exploitation of data-conditioned Hypernetworks in image-to-image tasks is substantial, considering their robust adaptability to diverse input images based on their content.